OpenStudy (anonymous):

Help! Please show work too. So confused. /: Solve the quadratic equation x2 - 4x + 57 = -5 by completing the square.

OpenStudy (anonymous):

completing square: \[x^{2} + bx = (x + (b/2))^{2} - (b/2)^{2}\]

OpenStudy (anonymous):

\[x^{2} - 4x = (x - 2)^{2} - 4\] were looking at: \[(x-2)^{2} - 4 + 57 +5 = 0\]

OpenStudy (anonymous):

let me know if i should solve for x (rearranging)